Find the number whoes product is 150 such that the one of the number is one more then four times the other number
Answer by ptaylor(2198)

Let x=one number
Then 4x+1=the other number
And we are told that:
x(4x+1)=150
4x^2+x=150 subtract 150 from each side
4x^2+x-150=0 quadratic in standard form
If we solve using the quadratic formula, we get:
x=(-1+-49)/8 or
x=-50/8=-25/4
and
x=48/8=6
Two solutions
CK
x=-25/4
4x+1=4(-25/4)+1=-24
(-25/4)(-24)=-25*-6=150
150=150
x=6
4x+1=25
6*25=150
150=150
Hope this helps--ptaylor